# Algebraic geometry of bubbling Kähler metrics

### v1

@@ 3.1. Deeper bubbles for klt degenerations
In the above setup, we further assume $\mathcal{X}$ descends to the family over $\Delta$ i.e., $\mathcal{X}=\mathcal{Y}$ for simplicity. Consider the projective system $(S,>)$ of isomorphism classes of $\varphi\colon\tilde{\mathcal{X}}\to\mathcal{X}$ which run over birational projective morphism such that $E_{x}$ is a divisorial part of $\varphi^{-1}(x)$ and $(\tilde{\mathcal{X}},\tilde{\mathcal{X}}_{0}+E_{x})$ is dlt and $\mathbb{Q}$-factorial, so that the dual intersection complex $\Delta(E_{x})$ is well-defined. (In case $x\in X$ is an isolated singularity, we can instead put stronger assumptions $\tilde{\mathcal{X}}$ is smooth and $\tilde{\mathcal{X}}_{0}+E_{x}$ is simple normal crossing for simplicity.) $(\varphi_{1}\colon\tilde{X}_{1}\to\mathcal{X})>(\varphi_{2}\colon\tilde{X}_{2}\to\mathcal{X})$ is simply defined for domination by a commuting birational morphism $\mathcal{X}_{1}\to\mathcal{X}_{2}$.
@@
    where $\varphi\in S$ and the domination “$>$” induces PL maps as in loc.cit. In particular, (14) is a dense inclusion while (15) is not.
@@
3.2. Towards bubbling for log canonical degeneration
@@
We end our paper by making some further discussions towards understanding of bubbling along worse degenerations than log terminal singularities, following our previous works ([Od22a, Od22b]). This subsection is independent of other (sub)sections. As a common setup of diverging or collapsing Ricci-flat Kähler metrics in the literature, one often consider the cases when the section hits a non-log-terminal, but strictly (semi-)log canonical point, like a non-smooth point at simple normal crossing degeneration. In such a situation, natural metric limits are expected to collapse. Therefore, to extend our approach, one can instead naturally ask the existence of minimal non-collapsing bubble (with infinite diameter) as defined in [Od22a, Definition 4.2] if it exists. Nevertheless the situation is more intricate (cf., e.g., [Od22a, 4.4], [HSVZ22, SRZ23]) and far beyond the main setup of this paper.
Here, as we expect it to play some roles in such general bubbling problem, we first review a simpler local version ([Od22b, Definition 2.5, Proposition 2.6]) of the CM degree minimization ([Od12, §4]) as follows. The following is a variant of weight function by [KS06, MN15, Tem14].
Definition 3.3 (Log canonical height).
For a projective flat generically K-trivial191919only means the canonical divisor of the generic fiber is numerically trivial family over a smooth proper curve $\mathcal{X}\to C\ni 0$ and a holomorphic section $\sigma\colon C\to\mathcal{X}$, where $\mathcal{X}$ is normal and $\mathbb{Q}$-factorial, we define
$h_{\rm lc}(\sigma):=\operatorname{deg}_{C}\hskip 2.84526pt\sigma^{*}(\mathcal{O}_{\mathcal{X}}(K_{\mathcal{X}/C}+\mathcal{X}_{0,{\rm red}})),$
and call it the log canonical height of $\sigma$. Here, $\mathcal{X}_{0,{\rm red}}$ refers to the reduced scheme structure on $\mathcal{X}_{0}$.
Though we work over a proper base $C$, note that it is only a matter of formal convenience to make sense of the degree. Indeed, along any vertical birational transforms of $\mathcal{X}$, one can always make sense of the difference even over non-proper $C$. Simply for this convenience, we keep the proper base setup. The following observation tells us the importance of log canonical minimal models.
Proposition 3.4.
For a flat proper family of log terminal projective K-trivial varieties $\mathcal{X}^{*}\to(C\setminus\{0\})$, we fix the meromorphic section $\sigma|_{C\setminus 0}$. Here, $C$ is a smooth proper curve. We assume $\mathcal{X}^{*}$ is $\mathbb{Q}$-factorial. Then among all fillings i.e., proper flat $\mathcal{X}\to C$ possibly after a finite base change of $C$ of degree $d$, we consider the quantity
$h_{lc}^{\nu}(\sigma):=\frac{h_{lc}(\sigma)}{d}.$
Then, $h_{lc}^{\nu}(\sigma)$ is minimized if and only if $\mathcal{X}$ around $\sigma(0)$ is a germ of some relative log canonical minimal model $(\mathcal{X},\mathcal{X}_{0,{\rm red}})$ over $C$.
@@
Firstly, if $(\mathcal{X},\mathcal{X}_{0,{\rm red}})$ is not log canonical around $\sigma(0)$, by the log canonical modification ([OX12]) $h_{\rm lc}(\sigma)$ would decrease by the negativity lemma. Therefore, we can and do assume $(\mathcal{X},\mathcal{X}_{0,{\rm red}})$ is at least log canonical locally around $\sigma(0)$. We replace $(\mathcal{X},\mathcal{X}_{0,{\rm red}})$ by a log canonical pair, keeping isomorphism around $\sigma(0)$ by [HX13], and run the relative minimal model program with ample $\mathbb{Q}$-Cartier scaling $\mathcal{H}$ to obtain the relative log canonical models of $(\mathcal{X},\mathcal{X}_{0,{\rm red}}+a\mathcal{H})$, by [HH20, Theorem 1.5]. We can assume $\sigma(0)$ is not contained in $\mathcal{H}$. If the corresponding divisorial contraction or log canonical flip is not locally isomorphism around $\sigma(0)$, then the log canonical height will decrease by the negativity lemma again. ∎
Recall that among the relative log canonical minimal models showed up above, there are dlt minimal models which are more special and simpler. Further, there is the technique of replacing the log canonical minimal model by its blow up which is dlt minimal model, in a log crepant manner, the so-called (log crepant) dlt modification or dlt blow up (cf., e.g., [Fuj11, KK10, OX12]). Thus, it sounds very natural to consider dlt minimal models to study these general bubbles, which led to the notion of “galaxy models” ([Od22b, §2]): connected locally ringed space which contains infinitely many bubbling log Calabi-Yau varieties densely.
Acknowledgement
The author heartily thanks S.Sun and C.Spotti for the helpful conversations over many years on this topic. He also thanks S.Boucksom, S.K.Donaldson, D.Halpern-Leistner, M.Jonsson, M.Mauri and T.Nghiem for helpful comments and discussions. During the research, the author was partially supported by Grant-in-Aid for Scientific Research (B) 23H01069, 21H00973, Grant-in-Aid for Scientific Research (A) 21H04429, 20H00112, and Fund for the Promotion of Joint International Research (Fostering Joint International Research) 23KK0249. Part of this work is done during the author’s stay at Paris, notably Institut de mathématiques de Jussieu - Paris Rive Gauche from March 2024 and Imperial College London in June 2024. The author received there above-mentioned benefitial comments, and nice pleasant working environments. We appreciate their kind hospitality.

### v2

@@ 3.1. Deeper bubbles for klt degenerations
In the above setup, we further assume $\mathcal{X}$ descends to the family over $\Delta$ for simplicity. Consider the projective system $(S,>)$ of isomorphism classes of $\varphi\colon\tilde{\mathcal{X}}\to\mathcal{X}$ which run over birational projective morphism such that $E_{x}$ is a divisorial part of $\varphi^{-1}(x)$ and $(\tilde{\mathcal{X}},\tilde{\mathcal{X}}_{0}+E_{x})$ is dlt and $\mathbb{Q}$-factorial, so that the dual intersection complex $\Delta(E_{x})$ is well-defined. (In case $x\in X$ is an isolated singularity, we can instead put stronger assumptions $\tilde{\mathcal{X}}$ is smooth and $\tilde{\mathcal{X}}_{0}+E_{x}$ is simple normal crossing for simplicity.) $(\varphi_{1}\colon\tilde{X}_{1}\to\mathcal{X})>(\varphi_{2}\colon\tilde{X}_{2}\to\mathcal{X})$ is simply defined for domination by a commuting birational morphism $\mathcal{X}_{1}\to\mathcal{X}_{2}$.
@@
    where $\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$ refers to the set of valuations of $\mathcal{O}_{\mathcal{X},x}$ with the center $x$ which maps $t$ to $1$, $\varphi\in S$ and the domination “$>$” induces PL maps as in loc.cit. In particular, (17) is a dense inclusion while (18) is not.
@@
3.2. Towards bubbling for semi-log canonical degeneration
@@
We make some further discussions towards understanding of bubbling along worse degenerations than log terminal singularities, following our previous works ([Od22a, Od22b]). This subsection is independent of other (sub)sections. As a common setup of diverging or collapsing Ricci-flat Kähler metrics in the literature, one often consider the cases when the section hits a non-log-terminal, but strictly (semi-)log canonical point, like a non-smooth point at simple normal crossing degeneration. In such a situation, natural metric limits are expected to collapse. Therefore, to extend our approach, one can instead naturally ask the existence of minimal non-collapsing bubble (with infinite diameter) as defined in [Od22a, Definition 4.2] if it exists. Nevertheless the situation is more intricate (cf., e.g., [Od22a, 4.4], [HSVZ22, SRZ23]) and far beyond the main setup of this paper.
Here, we fix such degeneration of polarized Calabi-Yau varieties over a punctured disk $\Delta^{*}$ to which we cannot fill in any log terminal polarized Calabi-Yau variety as the central fiber. For that, we define some subclass of possible irreducible components of degenerations, which we conjecture to form a large subset of minimal non-collapsing limits of the Calabi-Yau metrics. In this case, it sounds very natural to consider dlt minimal models to study these general bubbles, which led to the notion of “galaxy models” ([Od22b, §2]): connected locally ringed space which contains infinitely many bubbling log Calabi-Yau varieties densely. Then, we have been conjecturing that
Conjecture 3.3 ([Od22a, §4], [Od22b, §2]).
For any non-log-terminal semi-log-canonical (locally stable in the sense of Kollár) degeneration of polarized log terminal projective Calabi-Yau varieties $(\mathcal{X},\mathcal{L})\to\Delta$, we enhance it to (any) polarized galaxy model $\tilde{\mathcal{X}}\to\mathbbm{k}[[t^{\mathbb{Q}}]]$ and consider the klt locus of $\tilde{\mathcal{X}}^{\rm klt}_{0}$ its central fiber $\tilde{\mathcal{X}}_{0}$.
Each of its connected component appear as the minimal non-collapsing (in the sense of [Od22a]) pointed Gromov-Hausdorff limits (bubbling) of the Calabi-Yau metrics, for some sequence of base points $p_{t}$ in the general fibers $\mathcal{X}_{t}$.
In the case of families of polarized abelian varieties this is confirmed by explicit matrix calculations after Siegel [Od19] and further we know these galaxy models represent the non-archimedean Calabi-Yau metric as confirmed essentially in [Gub07, Gub10] (see [Od22b, §2]). Although the converse to the above conjecture is not true for non-flat cases, the author believes that the above bubbles forms the basic large subclass of minimal non-collapsing limits in general.
Acknowledgements   The author heartily thanks S.Sun and C.Spotti for the helpful conversations over many years on this topic. He also thanks H.Blum, S.Boucksom, M.deBorbon, S.K.Donaldson, D.Halpern-Leistner, M.Jonsson, Y.Kusakabe, M.Mauri and T.T.Nghiem for helpful comments and discussions. During the research, the author was partially supported by Grant-in-Aid for Scientific Research (B) 23H01069, 21H00973, Grant-in-Aid for Scientific Research (A) 21H04429, 20H00112, and Fund for the Promotion of Joint International Research (Fostering Joint International Research) 23KK0249. Part of this work is done during the author’s stay at Paris, notably Institut de mathématiques de Jussieu - Paris Rive Gauche from March 2024 and Imperial College London in June 2024. The author received there above-mentioned benefitial comments, and nice pleasant working environments. We appreciate their kind hospitality.
Appendix A Higher $\Theta$-stable reduction revisited
Recall that [Od24b, §3] introduces a higher rank generalization of the stable reduction avoiding $\Theta$-strata [HL14, AHLH23] and the construction in our Theorem 2.4, 2.6 and the properness of the moduli of Calabi-Yau cone in [Od24a] are partly its applications (in the irregular case). Since our original proof in [Od24b, §3] depends on sophisticated stack theoretic framework and techniques ([AHHLR22, HL14]), here we give more scheme theoretic reexplanation a priori conditionally (the conditions are often satisfied in practice e.g., for the above-mentioned applications).
First, we give scheme theoretic version of the higher $\Theta$-strata.
We work over any field $\mathbbm{k}$. Let us set $\mathcal{M}:=[M/G]$, where $M$ is a finite type $\mathbbm{k}$-scheme acted by an algebraic $\mathbbm{k}$-group $G$. We follow and partially use the basic notations of [Od24b]. For instance, $N\simeq\mathbb{Z}^{r}$ is a lattice, $T=\mathbb{G}_{m}\otimes N$ is an algebraic $\mathbbm{k}$-split torus, $\tau\subset N\otimes\mathbb{R}$ is a rational polyhedral cone and corresponding affine toric variety is denoted as $U_{\tau}$. In addition, here we denote $m\colon T\times T\to T$ is the multiplication and $a\colon G\times M\to M$ is the action.
Definition A.1 (after [Od24b, §3]).
We call a higher $\Theta$-strata on $[Z^{+}/G]\subset\mathcal{M}$ for $G$-invariant locally closed subscheme $Z^{+}(\subset M)$ of type $\tau\subset N\otimes\mathbb{R}$ is locally liftable if the corresponding $\mathcal{Z}^{+}\subset\underline{\rm Map}(\Theta_{\tau}=[U_{\tau}/T],[M/G])$ with ${\rm ev}={\rm ev}_{(1,1,\cdots,1)}\colon\mathcal{Z}^{+}\simeq[Z^{+}/G]$ is given by the following data:
  • a Zariski open covering $\{M^{\prime}\}$ of an open neighborhood of $Z^{+}$ in $M$.
  • For each $M^{\prime}$, we have an algebraic $\mathbbm{k}$-morphism $M^{\prime}\to{\rm Hom}_{\rm alg.\mathbbm{k}-grp}(T,G)$ (homomorphism family), which we encode by $\psi\colon T\times M^{\prime}\to G$, that satisfies the following:
    1. ($\mathrm{i}$)
      $a\circ(\psi\times{\rm id}_{M^{\prime}})\colon T\times M^{\prime}\to M$ gives $a^{\prime}\colon T\times M^{\prime}\to G\times M\to M$ whose image lies in $M^{\prime}$ and is a $T$-action (action-ness),
    2. ($\mathrm{ii}$)
      $a^{\prime}|_{(M^{\prime}\cap Z^{+})}$ extends to $U_{\tau}\times Z^{+}\to Z^{+}$.
    Here, ${\rm Hom}_{\rm alg.\mathbbm{k}-grp}(T,G)$ naturally means the algebraic $\mathbbm{k}$-scheme consists of homomorphisms of the algebraic $\mathbbm{k}$-groups, which naturally contains $G/Z(T)$ with the centralizer $Z(T)$ of $T$.
In practice, many examples of (higher) $\Theta$-strata satisfies the above local liftability condition and indeed we do not know any example which do not satisfies it. Under this a priori assumption, we can give a purely scheme-theoretic proof of the higher $\Theta$-stable reduction theorem of [Od24b]) as follows. We recall and formulate the statement in a simplified manner, as it is practically enough for above mentioned applications.
Theorem A.2 (cf., [Od24b, 3.8]).
Suppose a $\mathbbm{k}$-morphism $f\colon{\rm Spec}\mathbbm{k}[[t]]\to M$ satisfies that the closed point maps to a locally liftable (Definition A.1) higher $\Theta$-strata $Z^{+}$ and the generic point maps to outside $Z^{+}$. Replace $f$ by its base change $f\colon\mathbbm{k}[[t^{\frac{1}{d}}]]\to M$ for some $d\in\mathbb{Z}_{>0}$. Then, (as an elementary modification [AHLH23, §3.5]) there is a $\mathbbm{k}$-morphism $f^{\prime}\colon{\rm Spec}\mathbbm{k}[[t]]\to M\setminus Z^{+}$ with $f^{\prime}|_{\mathbbm{k}((t))}\in G(\mathbbm{k}((t)))\cdot f(\mathbbm{k}((t)))$.
The following arguments reproves Theorem 3.8 of [Od24b], but it does not depend on sophisticated results on algebraic stack theory, such as [AHHLR22], [HL14, 4.4.6] nor that of Białynicki-Birula type result ([HL14, 1.4.7, 1.4.8]). Nevertheless, the basic scheme-theoretic part of the proof of Theorem 3.8 of [Od24b] is still the same.
@@ Appendix A Higher $\Theta$-stable reduction revisited
We follow the proof of [Od24b, Theorem 3.8] but modify the arguments as follows. First, we take $(Z^{+})^{\prime}\subset M$ as Definition A.1 which contains $f(p)$ where $p$ denotes the closed point of ${\rm Spec}\mathbbm{k}[[t]]$. Set $a^{\prime}(p_{\tau},p)=:p^{\prime}$. Then we consider the $T$-action $a^{\prime}$ on a neighborhood of $p^{\prime}$ as $M^{\prime}\subset M$ and resolve the indeterminancy of the rational map $U_{\tau}(\supset T)\times\Delta\dashrightarrow M^{\prime}$ by applying Lemma 3.11 of [Od24b]. We recall what the claim of loc.cit implies briefly: after replacing $\tau$ to smaller enough subcone, it gives an appropriate rational polyhedral subcone of $\tau^{\prime}\subset\tau\times\mathbb{R}_{\geq 0}$ which projects isomorphically to $\tau^{\prime}$, set $\hat{\tau}^{\prime}:=\tau^{\prime}+\mathbb{R}_{\geq 0}$ and consider the corresponding affine toric variety $U_{\hat{\tau}^{\prime}}$, as well as its $1$-dimensional orbit $V(\tau^{\prime})(\simeq\mathbb{G}_{m})$ which corresponds to $\tau^{\prime}$. Then, $f$ extends to $U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]\to M$ so that $V(\tau^{\prime})\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]$ maps outside $Z^{+}$, due to loc.cit.
Further, taking a rational ray $\mathbb{R}_{\geq 0}(v,d)\subset\tau^{\prime}$ for a primitive integral vector $v\in\tau$ and $d\in\mathbb{Z}_{>0}$, we obtain ${\rm Spec}\mathbbm{k}[[t^{\frac{1}{d}}]]\to U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]$. Composing with the above $U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]\to M$, we obtain the required replacement $f^{\prime}$ as a desired modification. ∎
We note that in the proof of our Theorem 2.4 where we applied the original higher $\Theta$-stable reduction (Theorem 3.8 of [Od24b]), we proved the local liftability of the higher $\Theta$-strata in the above sense (Definition A.1), so that our above easier arguments A.2 suffices.
Now we assume $\mathbbm{k}$ is algebraically closed field of characteristic 00.
Then, we similarly show that the higher $\Theta$-strata which appears in the proof of properness of the moduli of Calabi-Yau cones is locally liftable in the sense of Definition A.1, so that the above easier version (Theorem A.2) suffices.
Theorem A.3 ([Od24a, 3.12]).
Fixing any positive integer $n$, $V$, Reeb vector field $\xi$, consider the moduli stack $\mathcal{M}_{V}^{\rm ss}(\xi)$ of $n$-dimensional K-semistable Fano cones of the minimized normalized volume $V$ and Reeb vector field $\xi\in N\otimes\mathbb{R}$. Then, $\mathcal{M}_{V}^{\rm ss}(\xi)$ is universally closed i.e., satisfies the existence part of the valuative criterion of properness.
We follow the arguments of §3.2 and §3.5 of op.cit. We write $T:=N\otimes\mathbb{G}_{m}$.
Proof.
We prove by the valuative criterion. Recall that $\mathcal{M}_{V}^{\rm ss}(\xi)$ is defined as a quotient stack $[B_{V}^{\rm ss}/G]$ in loc.cit, and $B_{V}^{\rm ss}$ can be enlarged to a larger $G$-acted algebraic $\mathbbm{k}$-scheme $B_{V}$ which is stratified as $B_{V}^{\rm ss}=B_{V}(0)\subset B_{V}(1)\subset\cdots\subset B_{V}(m^{\prime})=B_{V}$ by the value of the minimized normalized volumes again as in [Od24a, §3.2]. Thanks to the boundedness result [XZ24], now one can and do assume that each $B_{V}(i)$ contains all $n$-dimensional K-semistable Fano cone of the same volume density $V_{i}$, with $V=V_{0}>V_{1}>\cdots V_{m^{\prime}}$.
We take an arbitrary $\mathbbm{k}$-morphism ${\rm Spec}\mathbbm{k}[[t]]\to B_{V}$ such that the image $f(p)$ of the closed point $p$ lies in a connected component $Z^{+}$ of $B_{V}(i)\setminus B_{V}(i-1)$. (Recall that by the arguments in [Od24a], following [LX14], it follows that any $\mathbbm{k}$-morphism $\mathbbm{k}$-morphism ${\rm Spec}\mathbbm{k}((t))\to B_{V}^{\rm ss}$ can be extended to such for sufficiently large $B_{V}$ and $i$).
[…]
Then, if we restrict to $Z^{+}=B$, $\pi^{-1}(B)=:\mathcal{U}_{B}\to B$ has $(T\times)T_{B}$-action and canonical ideal sequences $I_{\lambda_{i}}\subset\mathcal{O}_{\mathcal{U}_{B}}$ for a discrete subset $\lambda_{i}\subset S(B)\subset\mathbb{R}_{\geq 0}$ with $0=\lambda_{0}<\lambda_{1}<\cdots$, which satisfies that $I_{\lambda_{i}}/I_{\lambda_{i+1}}$ is locally free $\mathcal{O}_{B}$-module of rank $r_{i}$ ([Che24, 1.1 (also §5)]). Here, $S(B)$ is the holomorphic spectrum (in the sense of [DS17]) corresponding to the $\pi$-fibers over $B$. Thanks to the fiberwise $T=N\otimes\mathbb{G}_{m}$-action, one can decompose $I_{\lambda_{i}}$ as $\oplus_{m\in M=N^{\vee}}I_{\lambda_{i},m}$ to the finite dimensional $T$-eigenspaces. Then, each $I_{\lambda_{i},m}/I_{\lambda_{i+1},m}$ is locally free $\mathcal{O}_{B}$-module of the finite rank $r_{i,m}$.
Now, consider the K-semistable degeneration of $[X]=f(p)\in B_{V}$ as $X\rightsquigarrow W$ and realize it as the $T_{B}$-action for some inclusion $T_{B}\subset G$ induced by the splitting of $\mathcal{O}_{X}\supset I_{\lambda_{1}}|_{X}\supset I_{\lambda_{2}}|_{X}\supset\cdots$. The limit $[W]$ still lies inside $B_{V}$ by the construction of $B_{V}$ mentioned above, as the minimized normalized volume remains the same as that of $X$ (see §2.1 and [XZ24]). We extend this inclusion $T_{B}\subset G$ as a homomorphism family locally around $[W]$ as follows. Since the $G$-action on $\mathcal{U}_{B}$ and the fiberwise $T$-action commutes, for each $\lambda_{i}$ and $m\in M^{\prime}$, we have a canonical $T_{B}(\times T)$-equivariant splitting of $\mathbbm{k}$-vector spaces $I_{\lambda_{i},m}|_{W}\twoheadrightarrow(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})|_{W}$ due to the complete reductivity of $T_{B}$. We extend it as follows. If we consider $I_{\lambda_{i},m}$ and $(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})$ as $\mathcal{O}_{B}$-modules, they are both locally free and inherit the $T_{B}$-action by the canonicity of $I_{\lambda_{i}}$s. Then, if we shrink $B$ to a small enough $T_{B}$-invariant open affine neighborhood of $[W]$ in $B$ by Sumihiro’s theorem, the $T_{B}$-equivariant vector bundles restricted to it are the direct sum of $T_{B}$-equivariant line bundles (cf., [IS15, 2.5], [Pay08, 2.1], [Kly89]).
By the same arguments as loc.cit, the $T_{B}$-equivariant surjection between $I_{\lambda_{i},m}\twoheadrightarrow(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})$ is $T_{B}$-equivariantly splittable by the complete reductivity of $T_{B}$ as $(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})\to I_{\lambda_{i},m}$ which corresponds to $T_{B}$-semiinvariant fiberwise coordinates on $\mathcal{U}_{B}=\pi^{-1}(B)$. We extend it to $T_{B}$-semiinvariant functions on $\pi^{-1}(B_{V}^{o})$ for a $T_{B}$-invariant open affine neighborhood $B_{V}^{o}$ of $[W]$ in $B_{V}$ which vanish at the 00-section($\simeq B_{V}^{o}$). If we collect them for all $m\in M^{\prime}$ and many enough $i$, and compare with the embedding of $\mathcal{U}\to\mathbb{A}^{l}\times B_{V}$ (for the same $M^{\prime}$), then this gives family of linear transformation $g(b)\in G\subset{\rm GL}(l)$ for each $b\in B_{V}^{o}$, which a priori non-trivially varies over $B_{V}^{o}$.
Then the family of homomorphism $((Z^{+})^{\prime}:=)B_{V}^{o}\to{\rm Hom}(T_{B},G)$ given by $b\mapsto g(b)^{-1}T_{B}g(b)$ ensures the local liftability of the higher $\Theta$-strata structure on $Z^{+}=B$, so that Theorem A.2 applies. By iterating the same procedure $i$-times, we obtain the desired modification ${\rm Spec}\mathbbm{k}[[t]]\to B_{V}^{\rm ss}$ of $f$, so that we conclude the assertion by the valuative criterion. ∎